OpenStudy (anonymous):

can anyone explain how as n--->infinity sin(6/n)/sin(3/n) approach zero. consider sequence convergent

OpenStudy (anonymous):

sin(6/n) / sin(3/n)

OpenStudy (kinggeorge):

Thanks @myko In that case, you know that the top goes to 0 and the bottom goes to 0, so you can use L'hopitals rule. \[\lim_{n \rightarrow \infty}{\sin(6/n) \over \sin(3/n)}={\cos(6/n) \over \cos(3/n)} \cdot {-6/n^2 \over -3/n^2}=\lim_{n\rightarrow \infty}1 \cdot {2n^2 \over n^2}=2\]

OpenStudy (kinggeorge):

Are you sure this is supposed to go to 0?

OpenStudy (anonymous):

yes it should be 2